OpenStudy (anonymous):

PLEASE HELP who wants a MEDAL Find the perimeter of a rectangle with l = 3.2 m and w = 0.5 m. A. 1.6 m B. 3.2 m C. 3.7 m D. 7.4 m

OpenStudy (disco619):

 Just add all the Sides.... Simple!!!

OpenStudy (disco619):

Perimeter = l + l + w + w = 3.2 + 3.2 + 0.5 + 0.5 = 3.2 x 2 + 0.5 x 2 = 6.4 + 1 = 7.4 Perimeter of Rectangle is = 7.4 m

OpenStudy (anonymous):

so it is D

OpenStudy (disco619):

Yes.

OpenStudy (anonymous):

thanks here is your medal

OpenStudy (disco619):

Your are Welcome.